# Detecting and removing bloated dependencies in CommonJS packages

### v2

@@ 7.2 Applications Debloating
Our work differs conceptually in both scope and methodology, addressing challenges specific to the dynamic and scripting nature of JavaScript. While static analysis is effective in statically typed languages like Java, JavaScript’s runtime behavior and the CommonJS module system complicate such approaches. This dynamic nature, combined with features like require()’s parameterized imports and runtime dependency generation, limits the applicability of static debloating techniques. To address these challenges, we adopt a dynamic, trace-based approach that monitors OS file system interactions during runtime, effectively capturing dependencies missed by static analysis. Our approach bridges a critical gap in dependency debloating research, offering specific solutions for CommonJS projects and providing insights into the unique dependency management challenges posed by the npm ecosystem. Meanwhile, our findings underscore the high accuracy and reliability of analyzing code usage at the OS file system level.
@@ 8 Conclusion
In this paper, we conduct a comprehensive study on bloated dependencies in Node.js packages that use the CommonJS module system. We propose a trace-based approach, implemented through a tool DepPrune, which monitors the runtime behavior of these packages. By tracking interactions between their dependencies and the OS file system, DepPrune effectively identifies both direct and indirect bloated dependencies. We evaluate DepPrune on a novel dataset of 91 real-world, well-tested CommonJS packages. Our analysis reveals that 50.6% of the runtime dependencies are bloated, underscoring the widespread prevalence of dependency bloat in CommonJS packages, particularly contributed by indirect dependencies. We also show the prevalence of direct bloated dependencies and demonstrate the significant benefits of removing them from the dependency tree. This highlights the cascading effect of direct bloated dependencies, where their removal often eliminates a substantial number of associated indirect dependencies. In addition, we conduct a comparative study with the state-of-the-art dynamic coverage-based analysis and static analysis techniques, for detecting CommonJS bloated dependencies. Our findings reveal that DepPrune outperforms these techniques by achieving higher accuracy, avoiding misclassification caused by static limitations, and addressing challenges posed by dynamic and non-linear execution patterns in JavaScript.
Our study highlights an opportunity to further enhance code quality by leveraging trace-based methods like DepPrune to improve the accuracy of dependency debloating in dynamic languages like CommonJS. In the future, we wish to investigate how dependency debloating can assist developers by reducing notification fatigue and enhancing overall development efficiency.

### v3

@@ 7.2 Applications Debloating
Our work differs conceptually in both scope and methodology, addressing challenges specific to the dynamic and scripting nature of JavaScript. While static analysis is effective in statically typed languages like Java, JavaScript’s runtime behavior and the CommonJS module system complicate such approaches. This dynamic nature, combined with features like require()’s parameterized imports and runtime dependency generation, limits the applicability of static debloating techniques. To address these challenges, we adopt a dynamic, trace-based approach that monitors OS file system interactions during runtime, effectively capturing dependencies missed by static analysis. Our approach bridges a critical gap in dependency debloating research, offering specific solutions for CommonJS projects and providing insights into the unique dependency management challenges posed by the npm ecosystem. We believe these contributions distinguish our work from prior research and address a critical gap in dependency debloating for JavaScript projects.
@@ 8 Conclusion
In this paper, we conduct a comprehensive study on bloated dependencies in Node.js packages that use the CommonJS module system. We propose a trace-based approach, implemented in a tool DepPrune, which monitors the runtime behavior of these packages. By tracking interactions between their dependencies and the OS file system, DepPrune effectively identifies both direct and indirect bloated dependencies. We evaluate DepPrune on a novel dataset of 91 real-world, well-tested CommonJS packages. Our analysis reveals that 50.6% of the runtime dependencies are bloated, underscoring the widespread prevalence of dependency bloat in CommonJS packages, particularly contributed by indirect dependencies. Meanwhile, we highlight the importance of bloated direct dependencies and the cascading benefits of removing them, in order to eliminate a substantial number of associated indirect dependencies. We further validate DepPrune through a comparative study with the state-of-the-art dynamic coverage-based analysis and static analysis techniques. Our findings reveal that DepPrune outperforms these techniques by achieving higher accuracy, avoiding the misclassifications due to the limitations of static analysis, and addressing challenges posed by dynamic and non-linear execution patterns in JavaScript.
Our study highlights an opportunity to further enhance code quality by leveraging trace-based methods like DepPrune to improve the accuracy of dependency debloating in dynamic languages like CommonJS. In the future, we wish to investigate how dependency debloating can assist developers by reducing notification fatigue, raising awareness about the nebraskan package maintainer (Munroe, 2020) and enhancing overall development efficiency.